9. A driver notices that her 1150 kg car slows down from 85 km/h to 65 km/h in about 6.0 s
on the level when it is in neutral. Approximately what power (Watts and hp) is needed to
keep the car travelling at constant speed of 75 km/h?
1
Expert's answer
2020-05-25T10:57:16-0400

9. A driver notices that her 1150 kg car slows down from 85 km/h to 65 km/h in about 6.0 s

on the level when it is in neutral. Approximately what power (Watts and hp) is needed to

keep the car travelling at constant speed of 75 km/h?


Given data

Mass of car is "m=1150 kg"

Initial velocity is "v_0=85 km\/h =23.6 m\/s"

Final velocity is "v=65 km\/h=18.06 m\/s"

The acceleration of the car is

"a=\frac{v-v_0}{t}=\frac{18.06-23.6}{6.0}=-0.924 m/s^2"

The force acting on the car is

"F=ma=(1150 kg)(-0.924 m\/s^2)=-1062.7 N"

The amount of power is needed to keep the car travelling at constant speed of 75 km/h is

"P=Fv=(-1062.7)(75(5\/18))=-2.214*10^4 W"